12y^1/3 divded by 4y1/2
Answer by stanbon(57250)

12y^1/3 divded by 4y1/2
----
= (12/4)(y^(1/3))/(y^(1/2))
-----
= 3*y^(1/3 - 1/2)
---
= 3*y^[(2-3)/6]
---
= 3*y^(-1/6)
----
= 3/y^(1/6)
===============
Cheers,
Stan H.